OpenStudy (anonymous):

Please help! 3. Solve: (y - -3) = -1/2 * (x - 4) -show work please-

OpenStudy (wolfboy):

what are you trying to say?

OpenStudy (anonymous):

what do you mean? @Wolfboy

OpenStudy (wolfboy):

that was at @iambatman

OpenStudy (wolfboy):

XD mwahahaha anyway back to the problem is it solve or spimplify

OpenStudy (anonymous):

solve it, . like my teacher gave us the answer but, we have to show how she got that answer, the problem is : Which equation represents a line with slope - 1/2 that passes through (-3, 4) ? And the answer is : 2y + x = 5 now i need to show work on how she got that.

OpenStudy (wolfboy):

well you could graph it.

OpenStudy (wolfboy):

Start at (-3,4) and use the slope to find the x and y intercept.

OpenStudy (wolfboy):

I believe that could be a helpful start

OpenStudy (wolfboy):

im having a brain fart can someone help? @iambatman @nincompoop

OpenStudy (anonymous):

Which question?

OpenStudy (anonymous):

Solve: (y - -3) = -1/2 * (x - 4)?

OpenStudy (wolfboy):

second one apparently

OpenStudy (wolfboy):

"Which equation represents a line with slope - 1/2 that passes through (-3, 4) ? And the answer is : 2y + x = 5 now i need to show work on how she got that"

OpenStudy (anonymous):

So you have slope and the x, y coordinates, just plug it in?

OpenStudy (anonymous):

y=mx+b

OpenStudy (wolfboy):

yea but the final answer is not in that form apparently so im confused

OpenStudy (wolfboy):

thats what i was thinking to

OpenStudy (vishweshshrimali5):

Okay @Mclovin21, I see you got very close to the correct answer.

OpenStudy (anonymous):

Ok so we have slope, m = -1/2, and you have coordinates (x,y) = (-3, 4). We want it in y = mx+b form 4=-1/2(-3)+b solve for b, then put it all together.

OpenStudy (vishweshshrimali5):

Good, now multiply 2 both sides. And please: \(\color{blue}{\text{Originally Posted by}}\) @Mclovin21 Please help! 3. Solve: \((y - \color{red}(-3\color{red})) = -1/2 * (x - 4)\) -show work please- \(\color{blue}{\text{End of Quote}}\)

OpenStudy (vishweshshrimali5):

Do not forget to add those brackets.

OpenStudy (wolfboy):

ok so i guess i was thinking correctly @iambatman thanks for clarifying

OpenStudy (anonymous):

Np

OpenStudy (vishweshshrimali5):

Now can you tell me @Mclovin21 , what will you get by multiplying 2 both sides.

OpenStudy (anonymous):

umm 3 ?

OpenStudy (vishweshshrimali5):

@iambatman and @Wolfboy , you can also use this equation: \((y-y_1) = m(x-x_1)\) where m is the slope and \((x_1,y_1)\) is the point lying on the line. @Mclovin21 , no let me show you.

OpenStudy (anonymous):

Yes, I know of the point - slope form, but I thought it was easier if we explained it this way, since wolf was already doing it.

OpenStudy (wolfboy):

yea I thought since it was the most simple it was the best route to take

OpenStudy (vishweshshrimali5):